# Large Language Models Might Not Care What You Are Saying: Prompt Format Beats Descriptions

### v1

@@ Abstract
With the help of in-context learning (ICL), large language models (LLMs) have achieved impressive performance across various tasks. However, the function of descriptive instructions during ICL remains under-explored. In this work, we propose an ensemble prompt framework to describe the selection criteria of multiple in-context examples, and preliminary experiments on machine translation (MT) across six translation directions confirm that this framework boosts ICL perfromance. But to our surprise, LLMs might not necessarily care what the descriptions actually say, and the performance gain is primarily caused by the ensemble format, since the framework could lead to improvement even with random descriptive nouns. We further apply this new ensemble prompt on a range of commonsense, math, logical reasoning and hallucination tasks with three LLMs and achieve promising results, suggesting again that designing a proper prompt format would be much more effective and efficient than paying effort into specific descriptions. Our code will be publicly available once this paper is published.
@@ 1 Introduction
In-context learning (ICL) boosts the performance of large language models (LLMs) across numerous natural language processing (NLP) tasks, where LLMs are presented with in-context examples containing both input and ground truth output (Brown et al. 2020; Dong et al. 2023). Many works have verified the vital role of in-context examples in ICL (Wang et al. 2023; Wei et al. 2023). However, Min et al. (2022) find that ground truth labels might not be the key to ICL performance on classification tasks.
@@
The selection of in-context examples has been proven significant to the performance of ICL (Rubin, Herzig, and Berant 2022) and there have been various works on in-context example selection (Agrawal et al. 2023; Li et al. 2023; Ye et al. 2023). Besides diverse approaches of selecting examples, no existing work has tried to explicitly tell LLMs in what way those specific examples are selected. We hypothesize that if LLMs are prompted with instructions describing the properties of selected in-context examples, they might learn better from these examples, since instruction following is one of LLMs’ most important qualities nowadays (Ouyang et al. 2022; Peng et al. 2023; Zhang et al. 2024).
@@
Recently, Tang, Wang, and Wu (2024) propose to prompt LLMs with examples selected based on both word-level and syntax-level criteria for machine translation (MT) for better ICL performance. This inspires us to tell LLMs where different in-context examples come from when examples are selected by multiple methods.
@@
In our experiments on MT, we first select in-context examples based on lexical and syntactic similarity for each test input separately. Then we combine both to construct the complete set of examples, with half word-level examples and half syntax-level examples. Further, we devise a novel ensemble prompt framework, adding example-level instructions to describe that the following examples are with similar words or similar syntax.
@@
Experimental results on MT demonstrate that adding such ensemble prompt framework does improve LLMs’ performance over conventional prompts. Meanwhile, we find that when the example-level descriptions do not correspond to the source of in-context examples (e.g., examples are said to be with similar syntax when they are actually selected based on word-level similarity) or are completely nonsense (e.g., the descriptive noun ”words” or ”syntax” is replaced with a random noun or the qualifier ”similar” is replaced with ”different”), LLMs still benefit from the prompt. These surprising results indicate that in fact LLMs might not care what the descriptions say and are more sensitive to the prompt format. In other words, a proper format can be much more effective than well-designed descriptions in ICL.
@@
To further verify the superiority of the ensemble framework, we present empirical evaluations on commonsense, math, logical reasoning and hallucination benchmarks (including nine datasets in total) across three small-scale LLMs (Alpaca, Llama2 and Mistral) and one large-scale LLM (GPT-3.5). The novel prompt framework is able to achieve promising results even with the descriptive nouns in the prompt being random nouns, further suggesting that a proper prompt format would be much more effective and efficient compared with laborious design of detailed and specific descriptions.
@@
  • We analyze the effect of prompt descriptions on ICL performance and find that LLMs might not necessarily care what users actually say in descriptions, while they are more sensitive to the prompt format. To the best of our knowledge, this is the first time that the effect of descriptions in prompts is specifically studied.
  • We present a simple yet effective prompt framework that is preliminarily proven feasible on MT through comprehensive experiments across six translation directions. Promising experimental results on three LLMs (Alpaca, Llama2 and Mistral) further verify the superiority of the novel framework on a range of commonsense, math, logical reasoning and hallucination tasks (including nine datasets in total).
2 Prompting LLMs for Machine Translation
Primarily, we focus on MT, a typical generation task that many works have studied ICL on it. Recently, various approaches of selecting in-context examples have been proposed for MT (Agrawal et al. 2023; Kumar et al. 2023; Tang, Wang, and Wu 2024). However, to the best of our knowledge, no previous work has tried to make LLMs aware of in what way those specific in-context examples are selected.
We assume that LLMs would perform better when they are told the reasons for selecting those examples. Tang, Wang, and Wu (2024) select examples based on a combination of word-level and syntax-level criteria, which inspires us to present an ensemble prompt to enable LLMs to clearly know the reasons behind example sampling. In addition, to have a comprehensive idea of whether LLMs really know what is said in the descriptions, we design some prompt variants that are less meaningful or completely nonsense.
2.1 In-context Example Selection for Machine Translation
For word-level examples, we simply select them using BM25 (Bassani 2023). For syntax-level examples, we use the simplified polynomial algorithm proposed by Tang, Wang, and Wu (2024) to convert dependency trees into polynomials and compute syntactic similarity based on the Manhattan distances (Craw 2017) between polynomial terms. For simplicity, we select syntax-level examples based on the Top-$k$ results of the polynomial algorithm instead of the coverage-based approach proposed by Tang, Wang, and Wu (2024). For brevity, we denote the syntax-level algorithm by ”polynomial”.
To combine examples from word-level and syntax-level criteria, we simply concatenate them. For example, the first half of in-context examples are selected by BM25 while the remaining half are selected by the polynomial algorithm.
@@ 2.2 A New Ensemble Prompt Framework
First of all, we use the most regular prompt without any example-level descriptions as a baseline (referred to as Vanilla), which is shown in Figure 1.
@@
In the template, ”Task-level Instruction” instructs the model to do the current task (MT here). ”Example $A_{i}$” and ”Example $B_{i}$” denote the $i$-th example from selection approach $A$ (e.g., BM25) and $B$ (e.g., polynomial) respectively, all containing both source language inputs and target language translations. ”Test Input” refers to the source language input of the test sample, which requires the LLM to translate it into the target language.
Then, we add example-level descriptions for examples from different selection approaches and explicitly instruct the LLM to translate the test input. This prompt framework is referred to as Ensemble and is shown in Figure 2.
Figure 2: Template and Alpaca’s example of Ensemble.
In the template, ”Noun $A$” and ”Noun $B$” describe the examples from selection $A$ and $B$ respectively. For example, the two nouns can be ”words” and ”syntax” to properly describe examples selected by BM25 and polynomial respectively. In this way, we can conveniently control the example-level instructions to tell the LLM why those in-context examples are used.
Figure 3: Main results on XGLM and Alpaca, showing the performance gain of different prompts over the vanilla prompt, averaged over all six translation directions. Each cluster presents the results of a selection of in-context examples and each bar in it presents the result of a prompt. ”Ens.”, ”W.”, ”Syn.”, ”Sem.”, ”Diff.”, ”Rand.”, ”Poly.” refer to ”Ensemble”, ”Word”, ”Syntax”, ”Semantics”, ”Different”, ”Random”, ”Polynomial” respectively.
@@ 2.3 Experimental Setup
Datasets
We perform evaluation on the devtest set of FLORES-101 (Goyal et al. 2022), which contains 1012 sentences with translations in 101 languages. We experiment between English and three common languages: German, French and Russian. We use Europarl (Koehn 2005) for German and French and ParaCrawl (Bañón et al. 2020) for Russian as example database, from which we select in-context examples. Detailed statistics are presented in Table 1.
Evaluation Metrics
We report COMET (Rei et al. 2020) scores from wmt20-comet-da  [^1] https://huggingface.co/Unbabel/wmt20-comet-da, which is considered a superior metric for MT today (Kocmi et al. 2021).
Language Models
@@
  • BM25: We retrieve the top-$4$ matching examples for each test input using BM25 (Bassani 2023).
  • Polynomial: It is rather time-consuming to retrieve examples from databases containing millions of data using the polynomial algorithm. Following Tang, Wang, and Wu (2024), we instead re-rank the top-$100$ examples retrieved by BM25 using polynomial and the top-$4$ are used as final in-context examples.
@@
  • Ensemble (Word + Syntax): The Ensemble prompt shown in Figure 2, with Noun $A$ and Noun $B$ being ”words” and ”syntax” respectively, which semantically corresponds to BM25 + Polynomial examples but is converse to Polynomial + BM25.
  • Ensemble (Syntax + Word): Noun $A$ and Noun $B$ are ”syntax” and ”words” respectively, which semantically matches Polynomial + BM25 examples but mismatches BM25 + Polynomial.
  • Different Ensemble (Word + Syntax): Noun $A$ and Noun $B$ are still ”words” and ”syntax” respectively but the qualifier ”similar” is replaced with ”different”. This prompt aims to find out whether LLMs know the meaning of ”different/similar” and care the semantics of descriptions.
  • Ensemble (Word + Semantics): Noun $A$ and Noun $B$ are ”words” and ”semantics” respectively.
  • Ensemble (Random + Random): For each test input, Noun $A$ and Noun $B$ are different random English nouns sampled using Wonderwords [^2] https://github.com/mrmaxguns/wonderwordsmodule. This aims to further explore LLMs’ understanding of descriptions.
@@ 2.4 Main Results
To give a quick view of LLMs’ performance on MT, Table 2 presents the COMET scores of baselines averaged over all six translation directions using Vanilla prompt.
@@
Main results are shown in Figure 3. For convenient comparison, we present the performance gain of different Ensemble prompts over Vanilla with different selections of in-context examples and the results are averaged over all six translation directions.
As can be seen from the results, those ”correct” prompts, exactly corresponding to the selection of in-context examples (e.g., Ensemble (Word + Syntax) with BM25 + Polynomial examples and Ensemble (Syntax + Word) with Polynomial + BM25 examples), do bring some help as expected. However, when the prompt does not correspond to the selection of examples (is not ”correct”), the performance improves as well and sometimes even more than those ”correct” cases. For example, on XGLM with BM25 + Polynomial examples, Ensemble (Syntax + Word) improves more than Ensemble (Word + Syntax), even though the former is completely reversed. On Alpaca with BM25 + Polynomial examples, Ensemble (Word + Semantics) improves more than Ensemble (Word + Syntax), albeit the examples with similar syntax do not necessarily bear similar semantics. More interestingly, Different Ensemble (Word + Syntax), telling the LLM that the in-context examples are with different properties, is able to beat ”correct” prompts sometimes (e.g., on XGLM with BM25 + Polynomial examples and Alpaca with Polynomial + BM25 examples).
Surprisingly, no matter how in-context examples are selected and whether the prompts are ”correct”, Ensemble prompts bring improvement in most cases. Even Ensemble (Random + Random), in which example-level descriptions are with random nouns and could be completely nonsense (like ”examples with similar nobody”), brings improvement in most cases, especially obtaining the most gain on Alpaca with Polynomial + BM25 examples compared with other prompts, correct or incorrect. These results indicate that LLMs might not really take the example-level descriptions into consideration during ICL. In other words, they might not necessarily care what users say in the descriptions.
@@ 2.5 Ablation Study
  • Single (Random): Organized based on Figure 2, but the second description with Noun $B$ is removed. There is only one example-level description above all the four examples, with Noun $A$ being a random noun.
  • Single (Example): Organized based on Figure 2, but the second description with Noun $B$ is removed. There is only one example-level description above all the four examples, being ”Examples:” only, without any further descriptions. This prompt only informs the LLM that the following four instances are examples and does not describe their properties.
  • Vanilla (Translate): Organized based on Figure 2, but both the two descriptions are removed. The only difference with Vanilla is the translation instruction ”Translate the following sentence:” before the test input. This prompt only informs the LLM to translate the test input and tells nothing about the in-context examples.
Figure 4: Ablation studies over the organization of the prompt, showing the performance gain of different prompts over Vanilla, averaged over all six translation directions. ”Rand.”, ”Poly.”, ”Ens.”, ”Sgl.”, ”V.”, ”Trans.” refer to ”Random”, ”Polynomial”, ”Ensemble”, ”Single”, ”Vanilla”, ”Translate” respectively.
@@
The ablation experiments demonstrate that in MT, our proposed Ensemble is a relatively superior prompt format, obtaining better results than other prompt variants.
@@ 2.6 Analysis via Attention Weights
From-To Attention Weight (×1e-4)
Ens. (W. + Syn.) Ens. (Rand. + Rand.)
EA - NA 69.03 60.63
EB - NA 22.49 23.28
EB - NB 16.42 22.93
T - NA 10.41 18.49
T - NB 10.89 18.24
Table 3: Results of attention weights on XGLM, comparing Ensemble (Word + Syntax) and Ensemble (Random + Random). ”EA”, ”EB”, ”NA”, ”NB” and ”T” denote examples from $A$ and $B$, Noun $A$ and Noun $B$, and the target position, respectively.
To have a better idea of the internal mechanism of LLMs when prompted with different prompts, we calculate the attention weights between different prompt parts. We focus on three components: in-context examples (from $A$ or $B$), the target position where the model starts to generate predictions (following Wang et al. (2023), we use the final token in the input) and the two descriptive nouns (Noun $A$ and Noun $B$). We obtain the attention weights averaged over all attention heads from the attention matrix in the last layer. Following Wang et al. (2023) and Zhao et al. (2021), we use the first subword if the noun is tokenized into multiple tokens. All the results are averaged over all six language directions.
Results on XGLM with BM25 + Polynomial examples are presented in Table 3. If the model really cares what the descriptions say, its attention to meaningful descriptive nouns should be much greater than those meaningless. However, the attention weights from examples to descriptive nouns of Ensemble (Word + Syntax) are not necessarily greater than those of Ensemble (Random + Random), indicating that the model does not care what the nouns actually are when managing in-context examples. The weights between the target position and the nouns of the former are even lower than the latter, showing that the model pays little attention to the actual meaning of the descriptions during generation. We conjecture that the relatively higher weights from the target position to random nouns might be due to the confusion of the model caused by those nonsense. To sum up, the attention weights further confirm our claims.

### v4

@@ Abstract
With the help of in-context learning (ICL), large language models (LLMs) have achieved impressive performance across various tasks. However, the function of descriptive instructions during ICL remains under-explored. In this work, we propose an ensemble prompt framework to describe the selection criteria of multiple in-context examples, and preliminary experiments on machine translation (MT) across six translation directions confirm that this framework boosts ICL performance. But to our surprise, LLMs might not care what the descriptions actually say, and the performance gain is primarily caused by the ensemble format, since it could lead to improvement even with random descriptive nouns. We further apply this new ensemble framework on a range of commonsense, math, logical reasoning and hallucination tasks with three LLMs and achieve promising results, suggesting again that designing a proper prompt format would be much more effective and efficient than paying effort into specific descriptions. Our code will be publicly available once this paper is published.
Large Language Models Might Not Care What You Are Saying:
Prompt Format Beats Descriptions
Chenming Tang  Zhixiang Wang  Hao Sun  Yunfang Wuthanks:    Corresponding author. National Key Laboratory for Multimedia Information Processing, Peking University MOE Key Laboratory of Computational Linguistics, Peking University School of Computer Science, Peking University {tangchenming, ekko}@stu.pku.edu.cn{2301213218, wuyf}@pku.edu.cn
@@ 1 Introduction
In-context learning (ICL) boosts the performance of large language models (LLMs) across numerous natural language processing (NLP) tasks, where LLMs are presented with in-context examples containing input and ground truth output Brown et al. (2020); Dong et al. (2023). Many works have verified the vital role of in-context examples in ICL Wang et al. (2023); Wei et al. (2023). However, Min et al. Min et al. (2022) find that ground truth labels might not be the key to ICL performance on classification tasks.
@@
The selection of in-context examples has been proven significant to the performance of ICL Rubin et al. (2022) and there have been various works on in-context example selection Agrawal et al. (2023); Li et al. (2023); Ye et al. (2023). Besides diverse approaches of selecting examples, no existing work has tried to explicitly tell LLMs in what way those specific examples are selected. We hypothesize that if LLMs are prompted with instructions describing the properties of selected in-context examples, they might learn better from these examples, since instruction following is one of LLMs’ most important qualities nowadays Ouyang et al. (2022); Peng et al. (2023); Zhang et al. (2024).
@@
Recently, Tang et al. (2024) prompt LLMs with examples selected based on both word-level and syntax-level criteria for machine translation (MT) for better ICL performance. This inspires us to tell LLMs where different in-context examples come from when they are selected by multiple methods.
Figure 1: Template and Alpaca’s example of Ensemble.
@@
In our experiments on MT, we first select in-context examples based on lexical and syntactic similarity for each test input separately. Then we combine both to construct the complete set of examples, with half word-level examples and half syntax-level examples. Further, we devise a novel ensemble prompt framework (as shown in the left part "Prompt Template" of Figure 1), adding example-level instructions to describe that the following examples are with similar words or similar syntax.
@@
Experimental results on MT demonstrate that adding such ensemble prompt framework does improve LLMs’ performance over conventional prompts. Meanwhile, we find that when the example-level descriptions do not correspond to the source of in-context examples or are completely nonsense, LLMs still benefit from the prompt. These surprising results indicate that in fact LLMs might not care what the descriptions say and are more sensitive to the prompt format. In other words, a proper format can be much more effective than well-designed descriptions in ICL.
@@
To further verify the superiority of the ensemble framework, we present empirical evaluations on commonsense, math, logical reasoning and hallucination benchmarks (including nine datasets in total) across three small-scale LLMs (Alpaca, Llama3 and Mistral) and one large-scale LLM (GPT-3.5). The novel prompt framework is able to achieve promising results even with the descriptive nouns in the prompt being random nouns, further suggesting that a proper prompt format would be much more effective and efficient compared with laborious design of detailed and specific descriptions.
There are a few studies very related to our work. Min et al. (2022) find that the labels of in-context examples do not need to be correct for classification tasks. Wei et al. (2023) find that larger language models successfully learn from in-context examples even when the labels are flipped or semantically-unrelated. Our work is different from the above in that we focus on the meaning of descriptions rather than labels in ICL and our finding is that the format of prompts is more important than carefully designed descriptions.
@@
  • For the first time, we specifically analyze the effect of prompt descriptions on ICL performance and find that LLMs might not care what users actually say in descriptions, while they are more sensitive to the prompt format.
  • We present a simple yet effective prompt framework that is proven feasible on MT through comprehensive experiments across six translation directions. Promising experimental results on three LLMs further verify the superiority of the novel framework on a range of commonsense, math, logical reasoning and hallucination tasks.
2 Prompting LLMs for MT
Primarily, we focus on MT, a typical generation task. Recently, various approaches of selecting in-context examples have been proposed for MT Agrawal et al. (2023); Kumar et al. (2023); Tang et al. (2024). However, no existing work has tried to make LLMs aware of in what way those specific in-context examples are selected.
We assume that LLMs would perform better when they are told the reasons for selecting those examples. Tang et al. Tang et al. (2024) select examples based on a combination of word-level and syntax-level criteria, which inspires us to present an ensemble prompt framework to make LLMs clearly know the reasons behind example selection. In addition, to have a comprehensive idea of whether LLMs really know what is said in the descriptions, we design some prompt variants that are less meaningful or completely nonsense.
2.1 In-context Example Selection for MT
For word-level examples, we simply select them using BM25 Bassani (2023). For syntax-level examples, we use the top-$k$ polynomial algorithm proposed by Tang et al. Tang et al. (2024) to convert dependency trees into polynomials and compute syntactic similarity based on the Manhattan distances Craw (2017) between polynomial terms. For brevity, we denote the syntax-level algorithm by "Polynomial".
To combine word-level and syntax-level examples, we simply concatenate them. For example, the first and the remaining half of examples are selected by BM25 and Polynomial respectively.
@@ 2.2 A New Ensemble Prompt Framework
First of all, we use the most regular prompt without any example-level descriptions as baseline (referred to as Vanilla), which is shown in Figure 2.
@@
In the template, "Task-level Instruction" instructs the model to do the current task (MT here). "Example $A_{i}$" and "Example $B_{i}$" denote the $i$-th example from selection approach $A$ (e.g., BM25) and $B$ (e.g., Polynomial) respectively, all containing both source language inputs and target language translations. "Test Input" refers to the source language input of the test sample, which requires the LLM to translate it into the target language.
Then, we add example-level descriptions for examples from different selection approaches and explicitly instruct the LLM to translate the test input. This prompt framework is referred to as Ensemble and is shown in Figure 1 as presented in Section 1. "Noun $A$" and "Noun $B$" describe the examples from selection $A$ and $B$ respectively. For example, the two nouns can be "words" and "syntax" to properly describe examples selected by BM25 and Polynomial respectively. In this way, we can conveniently control the example-level descriptions to tell the LLM why those examples are used.
Figure 3: Main results on XGLM and Alpaca, showing the performance gain of different prompts over the vanilla prompt, averaged over all six translation directions. Each cluster presents the results of a selection of in-context examples and each bar in it presents the result of a prompt. "Ens.", "W.", "Syn.", "Sem.", "Diff.", "Rand.", "Poly." refer to "Ensemble", "Word", "Syntax", "Semantics", "Different", "Random", "Polynomial" respectively.
@@ 2.3 Experimental Setup
2.3.1 Datasets
We perform evaluation on the devtest set of FLORES-101 Goyal et al. (2022), which contains 1012 sentences with translations in 101 languages. We experiment between English and three common languages: German, French and Russian. We use Europarl Koehn (2005) for German and French and ParaCrawl Bañón et al. (2020) for Russian as example database, from which we select in-context examples. Detailed statistics are in Table 1.
2.3.2 Evaluation Metrics
We report COMET Rei et al. (2020) scores from wmt20-comet-da  [^1] https://huggingface.co/Unbabel/wmt20-comet-da, which is considered a superior metric for MT today Kocmi et al. (2021).
2.3.3 Language Models
@@ 2.3.4 Example Selection
BM25: We retrieve the top-$4$ matching examples for each test input using BM25 Bassani (2023).
Polynomial: It is rather time-consuming to retrieve examples from databases containing millions of data using the Polynomial algorithm. Following Tang et al. Tang et al. (2024), we instead re-rank the top-$100$ examples retrieved by BM25 using Polynomial and the top-$4$ are used as final in-context examples.
@@ 2.3.5 Prompts
Ensemble (Word + Syntax): The Ensemble prompt shown in Figure 1, with Noun $A$ and Noun $B$ being "words" and "syntax" respectively, which semantically corresponds to BM25 + Polynomial examples but is converse to Polynomial + BM25.
Ensemble (Syntax + Word): Noun $A$ and Noun $B$ are "syntax" and "words" respectively, which semantically matches Polynomial + BM25 examples but mismatches BM25 + Polynomial.
Different Ensemble (Word + Syntax): Noun $A$ and Noun $B$ are still "words" and "syntax" respectively but the qualifier "similar" is replaced with "different". This prompt aims to find out whether LLMs know the meaning of "different/similar" and care the semantics of descriptions.
Ensemble (Word + Semantics): Noun $A$ and Noun $B$ are "words" and "semantics" respectively.
Ensemble (Random + Random): For each input, Noun $A$ and Noun $B$ are different random English nouns sampled using Wonderwords [^2] https://github.com/mrmaxguns/wonderwordsmodule, aiming to explore LLMs’ understanding of descriptions.
Detailed templates and examples of the above prompts are shown in Appendix A.1.
@@ 2.4 Main Results
To give a quick view of LLMs’ MT performance, Table 2 shows the COMET scores of Vanilla baselines averaged over six translation directions.
@@
Main results are shown in Figure 3. For convenient comparison, we present the performance gain of different Ensemble prompts over Vanilla with different selections of in-context examples and the results are averaged over six translation directions.
As can be seen from the results, those "correct" prompts, exactly corresponding to the selection of in-context examples (e.g., Ensemble (Word + Syntax) with BM25 + Polynomial examples and Ensemble (Syntax + Word) with Polynomial + BM25 examples), do bring some help as expected. However, when the prompt does not correspond to the selection of examples (is "incorrect"), the performance improves as well and sometimes even more than those "correct" cases. For example, on XGLM with BM25 + Polynomial examples, Ensemble (Syntax + Word) improves more than Ensemble (Word + Syntax), even though the former is completely reversed. On Alpaca with BM25 + Polynomial examples, Ensemble (Word + Semantics) improves more than Ensemble (Word + Syntax), albeit the examples with similar syntax do not necessarily bear similar semantics. More interestingly, Different Ensemble (Word + Syntax), telling the LLM that the in-context examples are with different properties, is able to beat "correct" prompts sometimes (e.g., on XGLM with BM25 + Polynomial examples and Alpaca with Polynomial + BM25 examples).
Surprisingly, no matter how in-context examples are selected and whether the prompts are "correct", Ensemble prompts bring improvement in most cases. Even Ensemble (Random + Random), in which example-level descriptions are with random nouns and could be completely nonsense (like "examples with similar nobody"), brings improvement in most cases, especially obtaining the most gain on Alpaca with Polynomial + BM25 examples compared with other prompts, correct or incorrect. These results indicate that LLMs might not really take the example-level descriptions into consideration during ICL. In other words, they might not necessarily care what users say in the descriptions.
@@ 2.5 Ablation Study
Single (Random): Organized based on Figure 1, but the second description is removed. There is only one example-level description above the four examples, where Noun $A$ is a random noun.
Single (Example): Organized based on Figure 1, but the second description is removed. There is only one example-level description above the four examples, being "Examples:" only, without any further descriptions. This prompt only informs the LLM that the following four instances are examples and does not describe their properties.
Vanilla (Translate): Organized based on Figure 1, but both the two descriptions are removed. The only difference with Vanilla is the translation instruction "Translate the following sentence:" before the test input. This prompt only informs the LLM to translate the test input and tells nothing about the in-context examples.
Detailed templates and examples of the above prompts are shown in Appendix A.2.
Figure 4: Ablation studies over the organization of the prompt, showing the performance gain of different prompts over Vanilla, averaged over all six translation directions. "Rand.", "Poly.", "Ens.", "Sgl.", "V.", "Trans." refer to "Random", "Polynomial", "Ensemble", "Single", "Vanilla", "Translate" respectively.
Figure 5: Attention weights (×1e-4) on XGLM of all 32 layers with BM25 + Polynomial examples. EWS and ERR denotes Ensemble (Word + Syntax) and Ensemble (Random + Random) respectively.
@@
Ablation experiments suggest that in MT, our proposed Ensemble is a relatively superior prompt format, performing better than other variants.
@@ 2.6 Analysis via Attention Weights
To have a better idea of the internal mechanism of LLMs when prompted with different prompts, we calculate the attention weights between different prompt components. We focus on three components: in-context examples (from $A$ or $B$, denoted by "Example-A" and "Example-B"), the target position (denoted by "Target") where the model starts to generate predictions (following Wang et al. Wang et al. (2023), we use the final token in the input) and the two descriptive nouns ("Noun-A" and "Noun-B"). We obtain the attention weights averaged over all attention heads from the attention matrix across all the layers. All the results are averaged over all six language directions.
Results comparing Ensemble (Word + Syntax) (EWS) and Ensemble (Random + Random) (ERR) on XGLM with BM25 + Polynomial examples are presented in Figure 5 (for results on Alpaca, refer to Appendix B). If the model really cares what the descriptions say, its attention to meaningful descriptive nouns (in EWS) should be much greater than those meaningless (in ERR). However, in most cases, EWS performs no higher than ERR, indicating that the model does not really care what the descriptive nouns actually are. "Target to Noun-A" is a special case, where EWS is high in shallow layers. But in deeper layers, EWS falls behind and ERR takes the lead. This shows that the model might pay more attention to the meaningful noun when understanding the context in shallow layers but gradually forgets it when it comes to generation in deeper layers. In a word, the attention weights further confirm our claim that LLMs do not really care what the descriptive nouns are in most cases.